Meeting Agenda

3 Feb 2004

In attendance:  J. Nelson (notetaker), J Frisch, S Smith, R Hall, J Rock, R Chestnut


New agenda items:

SLED FDBK through SCP slow feedback

     Ed Miller comes back 2/9.  Joe will try to cobble something together until then.


SIS waveforms archived every hour (plus on change)

     Janice will ask Marc if this is needed during the matlab proc loop running


Text string PV for shift taker name

     Judy will hook us up.


Visit of wisemen timeline

     Janice reported that the wiseguys are coming 4/26-27.  From on high, we’ve been given a mandate to hit the ground running when we come back from the structure install (~4/1)


Old agenda items:


List of phase 2a signals

Judy is working on the list.  After the meeting Janice met with Ron to spec out the processing loop for when we have 4 structures in the tunnel driven by 8-pack.  (We found one error in Joe’s doc – all structure in forward on the next line was called reflected.  Forward is correct.)



DIO problems

No more problems since database was fixed.  The error given was less than useless.


SISDTZ signal updates without archiving

Ron is working on this.


macnair module support of aes

Ron is working on this.  Maybe Joe and Linda can help?


Janice wants to mcaput values into proc loop fault counters.  Any foreseeable problems with this?

Ron says no problem.  He will let Janice know if any crop up.


Joe might need a new DIO channel for watching RF faults (diode trip isn't enough?)

Judy hooked us up right away.  Joe just got the cables to hook it up.  Testing will ensue.


Vac status into proc loop (T off SAM->VMEADC, need to talk to porter etal)
NO NEW STATUS – OPS (Janice will bug Marc again…)


epics display for plc vac analog display- analog trip statuses don’t turn red.  Maybe just not reboot?

Judy fixed this.  Yeah!


Button from Terri for hard gateway reset

     Janice is waiting to hear from Terri.


Summary PVs for plc panels? ongoing


Dac attens in MW instead of volts: (from last time) Ron found Steph’s clever solution and will look into what it takes to implement for 8pack. 

Janice will ask Marc if we still want/need this. NEED STATUS FROM RON – SEEMS LIKE WE’RE CANCELLING THIS REQUEST


PLC checkout




WE NEED ALAN HILL’S HELP or maybe Mark Larrus’s?


Lack of multiple archbr’s on one window – Ron Mackenzie fixed this.  yeah, thank you!



     Ron reported that Jingchen thinks it’s fixed.  It had a problem in its clock chip.  The SWE group is discussing re-arranging the processes that run on gtw4 vs gtw5.


Realm split

     This is ongoing.  Ops should expect just a few small interruptions in the future.